# A fast algorithm for 2D Rigidity Percolation

### v1

@@ Abstract
Rigidity Percolation is a crucial framework for describing rigidity transitions in amorphous systems. We present a new, efficient algorithm to study central-force Rigidity Percolation in two dimensions. This algorithm combines the Pebble Game algorithm, the Newman-Ziff approach to Connectivity Percolation, as well as novel rigorous results in rigidity theory, to exactly identify rigid clusters over the full bond concentration range, in a time that scales as $N^{1.02}$ for a system of $N$ nodes. Besides opening the way to accurate numerical studies of Rigidity Percolation, our work provides new insights on specific cluster merging mechanisms that distinguish it from the standard Connectivity Percolation problem.
@@ II.1 The Pebble Game
By construction, one run of the JH algorithm corresponds to a single value of the bond concentration $p$. The overall cost of the JH algorithm is determined by the number of pebble searches to be performed in a run –which depends on $N$– as well as the cost of each search –which also depends on $N$. JH reported a scaling of the algorithm that depends on $p$, scaling as $N^{1.15}$ near the critical point $p\approx p_{c}$, and linearly near full concentration. In practice this means that studying the phase diagram with the JH algorithm requires to run it for (as many as possible) different values of $p$, with a total computing time that depends on the number of values considered, most of which being typically concentrated near the critical point with cost $N^{1.15}$ each.
@@ V.4 Rigidification
To construct the rigid cluster resulting from rigidification, we start a Breadth First Search from $uv$ over roots of rigid clusters that could potentially coalesce into the forming one. To identify them, we introduce the Pivot Network $\mathcal{P}$, a graph whose nodes represent the rigid clusters existing in the system. If two rigid clusters share a pivot node in the physical system, they share an edge in the Pivot Network. Note that, from the “two-pivots rule” (Proposition 2), two distinct rigid clusters can share at most one pivot. In practice, the nodes of the Pivot Network correspond to the root bonds of the physical system, that uniquely identify each rigid cluster. The goal of the rigidification step is to identify the $k$ coalescing clusters/roots of Theorem 3, in order to construct $\mathcal{R}^{+}=\mathcal{R}_{1}\cup\mathcal{R}_{2}\cup\cdots\cup\mathcal{R}_{k}\cup\{uv\}$. Initially, $\mathcal{R}^{+}$ is composed of the new activated bond $uv$ only and both $u$ and $v$ are pivots of it. The BFS over the Pivot Network is performed by identifying the roots connected to $\mathcal{R}^{+}$ in the Pivot Network, and by checking the mutual rigidity of each of these roots with respect to $uv$. For each root $r$ found to be rigid with respect to $uv$, the corresponding rigid cluster is coalesced into $\mathcal{R}^{+}$, i.e., $\mathcal{R}^{+}\to\mathcal{R}^{+}\cup\mathcal{R}_{r}$. The pivots of each coalesced rigid cluster $\mathcal{R}_{r}$ become pivots of $\mathcal{R}^{+}$. The BFS then proceeds by further checking the new roots connected to $\mathcal{R}^{+}$ through these pivots. As shown in Fig. 3, this iterative procedure builds the rigid cluster resulting from rigidification by walking the paths in the Pivot Network that connect all the rigid clusters coalescing into $\mathcal{R}^{+}$. The BFS stops when a surrounding shell of floppy clusters has been identified.
@@ VI.1 Data structures
The system is represented by a standard adjacency list, the pebble graph is a $2N$ array where entries $2i$ and $12i+$ 1 are the pebble neighbors of node $i$. Negative values correspond to inactive pebble edges. We further define several other arrays of length $N$, the use of which is detailed in SM 4, including notably an array $\pi$, which stores the pivotal class of each node. The Pivot Network $\mathcal{P}$ is implemented as an array of sets with $M$ entries. C++ sets are balanced binary search trees that guarantee that no duplicates exist and, for a set with $n$ elements, they have $\mathcal{O}(\log n)$ cost of insertion, deletion and search  [^5] We preferred sets to unordered sets as we observed the former to perform better than the latter. Indeed, the complexity of basic operations on sets is guaranteed to be $\mathcal{O}(\log n)$. In contrast, while unordered sets offer $\mathcal{O}(1)$ complexity, their performance can degrade to $\mathcal{O}(n)$ due to rehashing of the underlying hash table.. Entries corresponding to root bonds contain the pivots of the corresponding rigid cluster. All other entries are empty sets. If the activation of $uv$ triggers a pivoting step, both $u$ and $v$ are inserted in the entry of bond $uv$ if their pivotal class becomes larger than one. Since in this case $u$ and $v$ become pivots of the other rigid cluster(s) they belong to, they must be inserted in the corresponding entries of $\mathcal{P}$ as well. We identify these rigid clusters by applying find operations on the active bonds adjacent to $uv$. The rigidification step is the more involved one, and we describe it in the next section. The Pivot Network is not affected by overconstraining.
@@ VI.2 Rigidification step
    Iterate over the bonds of $p$ For each bond, find its root $ij$. Once all the roots are processed as detailed below, go back to step 1 and process the next pivot.
@@ IX Conclusion
In this work we have presented novel theoretical results on two dimensional central-force Rigidity Percolation and have exploited them to derive an exact algorithm that computes the whole phase diagram of the model in a time that is almost linear in the system size. A C++ implementation of the algorithm is provided in [49]. Our algorithm pushes the size limitation encountered in previous works [5, 6, 30, 31], opening the possibility of accurately characterizing the RP universality class [41]. We have assessed the complexity intrinsic to the algorithm on the triangular lattice, where each node has exactly six neighbors. The rigidity transition, however, is of interest also on different topologies [26, 27]. Our algorithm can be straightforwardly implemented on any system where rigidity is ruled by Laman theorem. In particular, it can be readily used to analyze the rigidity of off-lattice configurations obtained from simulations of particle systems. On networks with a broad degree distribution, where the average degree might grow significantly with $N$ [51], the repeated application of the find operation would introduce a supplementary, $N$-dependent cost. In any other network as, e.g., Erdős–Rényi graphs we expect our algorithm to perform well. The actual performance might depend on fine details of the system such as the structure of the pebble graph (pebble searches might be more or less expensive than on the triangular lattice) and on the geometry of rigid clusters (number of pivots they have).
Our theoretical results provide new material to the long-standing question of the relationship between CP and RP [52, 23, 16], by showing explicitly that different cluster merging mechanisms are at play in the two problems. RP features notably non-local “rigidification” events that lead to the coalescence of several (rigid) clusters, a phenomenon absent in CP. In contrast with the standard Pebble Game, we construct the system in a dynamic way, making it straightforward to study accurately the frequency and spatial extension of rigidification events  [^6] While concluding this work we became aware of the preprint [57], that proposes a similar dynamic approach to RP. The algorithm has not been published, but is reported to perform slightly worse than the standard Pebble Game.. It would also be very interesting to generalize this approach to study rigidity loss [54], in particular large-scale fluidization events whereby the removal of a single bond makes a macroscopic rigid cluster collapse into smaller clusters connected by non-rigid hinges.
@@ X Acknowledgments
The authors are grateful to Lorenzo Rovigatti, Silvano Ferrari, Emanuela Bianchi and Claudio Castellano for insightful discussions and valuable comments on the manuscript. We especially thank Hugo Vanneuville and Vincent Beffara for their precious help in finalizing the proofs given in the SM.

### v2

@@ Abstract
Rigidity Percolation is a crucial framework for describing rigidity transitions in amorphous systems. We present a new, efficient algorithm to study central-force Rigidity Percolation in two dimensions. This algorithm combines the Pebble Game algorithm, the Newman-Ziff approach to Connectivity Percolation, as well as novel rigorous results in rigidity theory, to exactly identify rigid clusters over the full bond concentration range, in a time that scales as $N^{1.02}$ for a system of $N$ nodes. We perform extensive numerical simulations with systems larger than $500$ million nodes, far beyond the previous limitations. We obtain new, precise estimates for the critical exponents, $\nu=1.1694(8)$ and $D_{f}=1.8423(7)$, and locate the critical threshold at $p_{c}=0.6602741(4)$. Besides opening the way to further accurate numerical studies of Rigidity Percolation, our work provides new rigorous theoretical insights on specific cluster merging mechanisms that distinguish it from the standard Connectivity Percolation problem.
@@ II.1 The Pebble Game
By construction, one run of the JH algorithm corresponds to a single value of the bond concentration $p$. The overall cost of the JH algorithm is determined by the number of pebble searches to be performed in a run –which depends on NN– as well as the cost of each search –which also depends on $N$. JH reported a scaling of the algorithm that depends on $p$, scaling as $N^{1.15}$ near the critical point $p\approx p_{c}$, and linearly near full concentration. In practice this means that studying the phase diagram with the JH algorithm requires to run it for (as many as possible) different values of $p$, with a total computing time that depends on the number of values considered, most of which being typically concentrated near the critical point with cost $N^{1.15}$ each.
@@ V.4 Rigidification
To construct the rigid cluster resulting from rigidification, we start a Breadth First Search from uvuv over roots of rigid clusters that could potentially coalesce into the forming one. To identify them, we introduce the Pivot Network $\mathcal{P}$, a graph whose nodes represent the rigid clusters existing in the system. If two rigid clusters share a pivot node in the physical system, they share an edge in the Pivot Network. Note that, from the “two-pivots rule” (Proposition 2), two distinct rigid clusters can share at most one pivot. In practice, the nodes of the Pivot Network correspond to the root bonds of the physical system, that uniquely identify each rigid cluster. The goal of the rigidification step is to identify the $k$ coalescing clusters/roots of Theorem 3, in order to construct $\mathcal{R}^{+}=\mathcal{R}_{1}\cup\mathcal{R}_{2}\cup\cdots\cup\mathcal{R}_{k}\cup\{uv\}$. Initially, $\mathcal{R}^{+}$ is composed of the new activated bond $uv$ only and both $u$ and $v$ are pivots of it. The BFS over the Pivot Network is performed by identifying the roots connected to $\mathcal{R}^{+}$ in the Pivot Network, and by checking the mutual rigidity of each of these roots with respect to $uv$. For each root $r$ found to be rigid with respect to $uv$, the corresponding rigid cluster is coalesced into $\mathcal{R}^{+}$, i.e., $\mathcal{R}^{+}\to\mathcal{R}^{+}\cup\mathcal{R}_{r}$. The pivots of each coalesced rigid cluster $\mathcal{R}_{r}$ become pivots of $\mathcal{R}^{+}$. The BFS then proceeds by further checking the new roots connected to $\mathcal{R}^{+}$ through these pivots. As shown in Fig. 3, this iterative procedure builds the rigid cluster resulting from rigidification by walking the paths in the Pivot Network that connect all the rigid clusters coalescing into $\mathcal{R}^{+}$. The BFS stops when a surrounding shell of floppy clusters has been identified.
@@ VI.1 Data structures
The system is represented by a standard adjacency list, the pebble graph is a $2N$ array where entries $2i$ and $2i+1$ are the pebble neighbors of node $i$. Negative values correspond to inactive pebble edges. We further define several other arrays of length $N$, the use of which is detailed in SM 4, including notably an array $\pi$, which stores the pivotal class of each node. The Pivot Network $\mathcal{P}$ is implemented as an array of sets with $M$ entries. C++ sets are balanced binary search trees that guarantee that no duplicates exist and, for a set with $n$ elements, they have $\mathcal{O}(\log n)$ cost of insertion, deletion and search  [^5] We preferred sets to unordered sets as we observed the former to perform better than the latter. Indeed, the complexity of basic operations on sets is guaranteed to be $\mathcal{O}(\log n)$. In contrast, while unordered sets offer $\mathcal{O}(1)$ complexity, their performance can degrade to $\mathcal{O}(n)$ due to rehashing of the underlying hash table.. Entries corresponding to root bonds contain the pivots of the corresponding rigid cluster. All other entries are empty sets. If the activation of $uv$ triggers a pivoting step, both $u$ and $v$ are inserted in the entry of bond $uv$ if their pivotal class becomes larger than one. Since in this case $u$ and $v$ become pivots of the other rigid cluster(s) they belong to, they must be inserted in the corresponding entries of $\mathcal{P}$ as well. We identify these rigid clusters by applying find operations on the active bonds adjacent to $uv$. The rigidification step is the more involved one, and we describe it in the next section. The Pivot Network is not affected by overconstraining.
@@ VI.2 Rigidification step
    Iterate over the bonds of $p$ . For each bond, find its root $ij$. Once all the roots are processed as detailed below, go back to step 1 and process the next pivot.
@@ VIII.1 Performance of the algorithm
VIII.2 Estimation of the RP critical exponents and critical threshold
We apply our algorithm to the determination of the main critical exponents – the correlation length exponent $\nu$, the fractal dimension $D_{f}$, the susceptibility exponent $\gamma$ and the order parameter exponent $\beta$. We also determine the critical point $p_{c}$. We perform simulations for system sizes ranging from $L=2^{6}=64$ to $L=2^{14.5}=23170$, and choose the numbers of samples so as to activate at least $7\cdot 10^{11}$ bonds per value of $L$ (see SM 6). For each run of the algorithm, and for both the connectivity and rigid clusters, we store: the number of bonds $m_{x}$ and $m_{xy}$ at which wrapping happens for the first time, respectively along the horizontal and along both horizontal and vertical directions; the size $S_{xy}$ of the $xy$-wrapping cluster right at wrapping; the maximum value $\chi_{\rm max}$ of the susceptibility.
We then perform scaling analysis to determine $D_{f}$, $\nu$, $\gamma/\nu$ and $p_{c}$, for both the RP and CP transitions, the latter being used to benchmark our scaling analysis. The values of all exponents are reported in Table 1, giving a comparison of the CP and RP universality classes. These values are validated by the excellent collapse of the rescaled scaling functions (order parameter, susceptibility and wrapping probability) shown in Fig. 7. Overall, our results strengthen the long-standing hypothesis [5] that 2d CP and central-force RP belong to different universality classes.
VIII.2.1 Fractal dimension $D_{f}$
The fractal dimension $D_{f}$ is estimated from the (rescaled) size of the wrapping cluster $S_{xy}/(3L^{2})$, expected to scale as $S_{xy}\sim L^{D_{f}-2}$ [1]. For CP, a simple power-law fit gives an estimate in excellent agreement with the exact value $91/48$, reported in Figure 6 (a). For RP, deviations to scaling are present at small sizes, so that we choose to fit $S_{xy}^{\rm RP}$ as a power-law, but only for sufficiently large $L$. The number of datapoints removed is chosen as the one maximizing the goodness of fit, and indicated in Fig. 6. Note that we also tried to fit the data of $S_{xy}$ (and similarly for $\Delta$, $\chi_{\rm max}$ and $p_{c}$) to the form $\mathcal{O}(L)=A\,L^{y}\left(1+B/L^{w}\right)$, which incorporates corrections to scaling. For all observables, such fits lead values of the exponent $y$ consistent with those obtained by removing the smallest sizes from a power-law fit.
VIII.2.2 Susceptibility exponent $\gamma$
The maximum of the susceptibility is known to scale as $\chi_{max}\sim L^{\gamma/\nu}$ [1], from which we estimate the exponent $\gamma/\nu$. As for the size of the wrapping cluster $S_{xy}$ (cf. section VIII.2.1), the data for CP is perfectly fitted by a power-law, while RP shows significant deviations to scaling at small sizes, see Figure 6 (b). We adopt the same strategy as in section VIII.2.1 and fit only data with sufficiently large $L>L_{\min}$, the value of $L_{\rm min}$ being chosen as maximizing the goodness of fit. The susceptibility exponent $\gamma^{\rm RP}$ is then computed using the value of $\nu^{\rm RP}$ obtained independently in section VIII.2.3, and reported in Table 1.
VIII.2.3 Correlation length exponent $\nu$
The measurements of $m_{x}$ and $m_{xy}$ allows to compute the wrapping probabilities $R_{x}$ (along the horizontal axis), $R_{xy}$ (along both axes), $R_{e}$ (along either axis) and $R_{1}$ (along one axis), see [44] and Fig. S6. For each wrapping probability, we compute the width of the transition region $\Delta(L)$, as explained in SM 6 A. For each wrapping type, $\Delta(L)$ is expected to scale as $\Delta(L)\sim L^{-1/\nu}$ [1].
For CP, simple power-law fits of $\Delta(L)$ yield exponents in excellent agreement with the exact value $\nu^{\rm CP}=4/3$, for all four wrapping types. For RP, small deviations to scaling are visible in the residual plots, especially for $R_{x}$ and $R_{xy}$, while they appear suppressed in $R_{e}$ and $R_{1}$. We adopt the same strategy as in section VIII.2.1 and fit the data for $\Delta(L)$ as power-laws, discarding the smallest sizes as indicated in Figure 6 (c). The corresponding values of $1/\nu^{\rm RP}$ are reported in Figure 6 (c). Note that not all four wrapping probabilities are independent [40]. We find that the estimates from $R_{x}$ and $R_{xy}$ lead to the best collapse for all four wrapping probabilities, and therefore compute the final value of $\nu^{\rm RP}$, reported in Table 1, as the average of $\nu^{\rm RP}_{x}$ and $\nu^{\rm RP}_{xy}$.
Figure 6: Scaling analysis for RP. (a), (b) Size $S_{xy}/(3L^{2})$ of the wrapping cluster, and maximum $\chi_{\rm max}$ of the susceptibility, both rescaled by the total number of bonds, as function of $L$, shown for CP and RP. The lines are power-law fits of the data marked with filled symbols. (c) Widths $\Delta(L)$ of the transition region for each wrapping type x,xy,e, 1x,\,xy,\,e,\,1, as functions of the system size $L$. The slopes of the solid lines are reported in the legend and are obtained from power-law fits of the data marked with filled symbols. For the sake of readability, data from $R_{xy}$, $R_{e}$ and $R_{1}$ have been shifted by a factor 2, 4 and 6 respectively. (d) Finite-size critical thresholds $p_{c}/\Delta$, for each wrapping type, as functions of $\Delta^{-1}$. The slopes of the solid lines, reported in the legend, give the corresponding values of $p_{c}^{\rm RP}$. Data has been shifted and shown in log-log scale to ease readability.
VIII.2.4 Critical threshold $p_{c}$
The wrapping probabilities $R_{i}$ allow to further compute, for each wrapping type, the size-dependent transition thresholds $p_{c}(L)$, namely the bond concentration at which the emergence of a wrapping cluster is expected on a finite lattice of size $L$ (cf. SM 6 A). In general we expect $p_{c}(L)-p_{c}\propto L^{-1/\nu}\propto\Delta(L)$, where $p_{c}=p_{c}(L\to\infty)$. For RP, pronounced deviations to scaling are visible at small sizes, namely $p_{c}^{\rm RP}(\Delta)=p_{c}^{\rm RP}+A\,\Delta^{\rm RP}\left(1+B\,\Delta^{w}\right)$ for some positive exponent $w$. We consider the quantity $p_{c}^{\rm RP}/\Delta^{\rm RP}$ and fit
$y(x)\equiv\frac{p_{c}^{\rm RP}(\Delta^{\rm RP})}{\Delta^{\rm RP}}=p_{c}\;x+A$ (3)
for $L>L_{\rm min}$ sufficiently large, namely when $x=1/\Delta^{\rm RP}$ is sufficiently large that the correction term $\propto x^{-w}$ becomes negligible. For each wrapping type the value of $L_{\rm min}$ is chosen to maximize the goodness of fit. All four wrapping types yield excellent fits, as shown in Fig. 6 (d). Our final estimate of $p_{c}^{\rm RP}$ reported in Table 1 is the average of the estimates obtained from $R_{x}$ and $R_{xy}$, that have the lowest uncertainties.
Figure 7: Scaling functions for RP (main plots), and their collapse (insets) when rescaled using the values of the critical threshold and exponents in Table 1. (a) Order parameter $P_{\infty}$ (b) Susceptibility $\chi$ (c) Wrapping probability $R_{xy}$.
CP RP
$D_{f}$ $91/48\approx 1.895833$ $1.8423(7)$
$\nu$ $4/3\approx 1.333333$ $1.1694(8)$
$\gamma$ $43/18\approx 2.388889$ $1.928(3)$
$\beta$ $5/36\approx 0.138889$ $0.1844(8)$
$p_{c}$ $2\sin(\pi/18)\approx 0.347296$ $0.6602741(4)$
Table 1: CP and RP universality classes, and critical thresholds. The value of $\beta$ is computed using hyperscaling as $\beta=\nu(2-D_{f})$.
@@ IX Conclusion
In this work we have presented novel theoretical results on two dimensional central-force Rigidity Percolation and have exploited them to derive an exact algorithm that computes the whole phase diagram of the model in a time that is almost linear in the system size. A C++ implementation of the algorithm is provided in [49]. We have assessed the complexity intrinsic to the algorithm on the triangular lattice, where each node has exactly six neighbors. The rigidity transition, however, is of interest also on different topologies [26, 27]. Our algorithm can be straightforwardly implemented on any system where rigidity is ruled by Laman theorem. In particular, it can be readily used to analyze the rigidity of off-lattice configurations obtained from simulations of particle systems. On networks with a broad degree distribution, where the average degree might grow significantly with $N$ [51], the repeated application of the find operation would introduce a supplementary, $N$-dependent cost. In any other network as, e.g., Erdős–Rényi graphs we expect our algorithm to perform well. The actual performance might depend on fine details of the system such as the structure of the pebble graph (pebble searches might be more or less expensive than on the triangular lattice) and on the geometry of rigid clusters (number of pivots they have).
Our algorithm pushes the size limitation encountered in previous works [5, 6, 30, 31], allowing us to characterize accurately the universality class: we have performed scaling analysis on systems up to 536 848 900536\,848\,900 nodes, and provide new estimates of the main critical exponents, $\nu$, $\beta$, $D_{f}$ and $\gamma$, as well as the critical threshold $p_{c}$. The consistency of these values is validated by the excellent collapse of the rescaled scaling functions, shown in Fig. 7.
On the theoretical side, our results provide new material to the long-standing question of the relationship between CP and RP [52, 23, 16], by showing explicitly that different cluster merging mechanisms are at play in the two problems. RP features notably non-local “rigidification” events that lead to the coalescence of several (rigid) clusters, a phenomenon absent in CP. In contrast with the standard Pebble Game, we construct the system in a dynamic way, making it straightforward to study accurately the frequency and spatial extension of rigidification events  [^6] While concluding the first version of this work we became aware of the preprint [61], that proposes a similar dynamic approach to RP. The algorithm has not been published, but is reported to perform slightly worse than the standard Pebble Game.. It would also be very interesting to generalize this approach to study rigidity loss [54], in particular large-scale fluidization events whereby the removal of a single bond makes a macroscopic rigid cluster collapse into smaller clusters connected by non-rigid hinges.
@@
Note added to the second version: While finalizing the second version of this work –which completes our original study by determining the critical exponents, we became aware of the preprint [57], which also reports new estimates for the RP critical exponents. To the best of our knowledge (the code not yet being publicly available), the algorithm described in [57] incorporates several strategies that we introduced in the first version of our work, but the system sizes considered in their analysis do not exceed $L=8192$. We observe that the reported values of the critical threshold and the correlation length exponent are in good agreement with our results. The larger value of the fractal dimension, $D_{f}=1.850(2)$, may be explained by the fact that it is obtained from a different set of observables, wrapping not being implemented in their algorithm.
@@ X Acknowledgments
The authors are grateful to Lorenzo Rovigatti, Silvano Ferrari, Emanuela Bianchi and Claudio Castellano for insightful discussions and valuable comments on the manuscript. We especially thank Hugo Vanneuville and Vincent Beffara for their precious help in finalizing the proofs given in the SM. Finally, we thank the organizers of the 2024 Network Dynamics workshop in Les Houches, where this project started.